OpenStudy (anonymous):

If f(x) = 5x + 3 and g(x) = 2x + 3, what is f(g(x)) ?

Elsa213 (elsa213):

f(g(x)) = f(2x+1) f(2x+1) = 5(2x+1) - 3 Just distribute now 10x+5-3 f(g(x)) = 10x+2

OpenStudy (chrisisdope1):

f(x)= 5x-3 f(g(x)) = 5(2x+1) -3 = 10x +5 - 3 = 10x +2

OpenStudy (anonymous):

A 10x + 21 B 10x + 18 C 10x + 15 D 10x + 12

OpenStudy (anonymous):

b is right answer

OpenStudy (chrisisdope1):

Yea He's Right

OpenStudy (anonymous):

f(g(x)) = f(2x+3) f(2x+1) = 5(2x+3) - 3 = 10x+15+3 = 10x+18

OpenStudy (anonymous):

thank yall so much!

OpenStudy (anonymous):

welcome